OpenStudy (anonymous):

if f(x)=4x^2 +7 and g(x)=5x+2. Find g(f(x))

OpenStudy (anonymous):

\[g(f(x))=g(5x+2)=4(5x+2)^2+7\] then some algebra to clean it up

OpenStudy (anonymous):

\[If f(x)=4x^2 +7 and g(x)=5x+2\]